OpenStudy (babynini):

Convert polar coordinates (8, 5pi/4) to rectangular coordinates.

OpenStudy (babynini):

I know r^2 = sqroot (x^2+y^2) but not sure how to do that with this one..

OpenStudy (babynini):

@satellite73

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

you do not have a curve, just a point use \[x=r\cos(\theta), y = r\sin(\theta)\]

OpenStudy (babynini):

oou

OpenStudy (babynini):

Give me a moment to calculate that :)

OpenStudy (babynini):

so they both equal \[-\frac{ 8\sqrt{2} }{ 2 }\]

OpenStudy (babynini):

which can be simplified to \[-\frac{ 4\sqrt{2} }{ 2 }\] right?

OpenStudy (babynini):

wait..no -4 sqroot2